OpenStudy (anonymous):

In need of serious help, with Solving Systems of Linear Equations by Substitution. the equations are 4x+y=22 5x-2y=8 PLEASE SOMEONE HELP ME OUT!!!

OpenStudy (anonymous):

you wanna find x and y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have to know either there's a Solution, There are infinitely many solutions or No Solutions

OpenStudy (anonymous):

f\[2x4x+y=22\] \[5x-2y=8\]

OpenStudy (anonymous):

from 2 x first equetion we have: 8x+2y=44 5x-2y=8

OpenStudy (anonymous):

then sum two equation so we have: 13x=52 then x=52/13

OpenStudy (anonymous):

I'm so lost now

OpenStudy (anonymous):

with replace x4 into one of two equation we have: 5*(4)-2y=8 then we have find y=6

OpenStudy (anonymous):

it's ok?